### v2

# Security of AI Agents

Abstract
The study and development of AI agents have been boosted by large language models. AI agents can function as intelligent assistants and complete tasks on behalf of their users with access to tools and the ability to execute commands in their environments, Through studying and experiencing the workflow of typical AI agents, we have raised several concerns regarding their security. These potential vulnerabilities are not addressed by the frameworks used to build the agents, nor by research aimed at improving the agents. In this paper, we identify and describe these vulnerabilities in detail from a system security perspective, emphasizing their causes and severe effects. Furthermore, we introduce defense mechanisms corresponding to each vulnerability with meticulous design and experiments to evaluate their viability. Altogether, this paper contextualizes the security issues in the current development of AI agents and delineates methods to make AI agents safer and more reliable.
1 Introduction
AI agents are robots in cyberspace, executing tasks on behalf of their users. To understand their user’s command, they send the input prompts as requests to foundation AI models, such as large language models (LLMs). The responses generated by the model may contain the final actions of the agent or further instructions. To execute the actions, the agent invokes tools, which may run local computations or send requests to remote hosts, such as querying search engines. The tools output results and feedback to the AI model for the next round of actions. By invoking tools, AI agents are granted the ability to interact with the real world. Since AI agents depend on their AI model to understand user input and environment feedback and generate actions to use tools, we say that the AI model is the backbone of the agent. We summarize the basic architecture of LLM-based AI agents in Figure 1. Traditional agents operate on pre-defined rules [castelfranchi1998modelling, wilkins2014practical] or reinforcement learning models [isbell2001social], making them hard to generalize to new tasks and different tools. LLM-based AI agents, on the contrary, can be practical in various tasks benefiting from enormous pre-training knowledge and the ability to read tool documentation as additional prompts. We use the term AI agent to denote all LLM-based agents in this paper.
Over the years, AI agents have showcased their outstanding performance on tasks including but not limited to writing shell scripts to interact with operating systems, querying databases, shopping and browsing on the web, and playing video games [yao2022webshop, park2023generative, zhou2023webarena, liu2023agentbench]. Agent-based personal assistants are also gaining popularity in society, for example, Microsoft [semantic_kernel, copilot] and HuggingFace [hfagents]. Overall, AI agents are a promising step forward toward the goal of robots with artificial general intelligence (AGI).
Figure 1: Overview of LLM-based AI agent.
Despite their popularity, existing research and development of AI agents failed to take into account their potential vulnerabilities. In traditional computing systems, security is guarded by three properties: confidentiality, integrity, and availability. Each of these properties faces unique challenges in AI agent usage.
Confidentiality is often managed by model-based access control policies, which abstract the system components and users into subjects, objects, and rights [bishop2004introduction]. However, these principles face significant challenges when applied to LLM-based systems due to the nature of LLMs to memorize [mireshghallah2022memorization, carlini2023quantifying, tirumala2022memorization] and compress [deletang2023language] training data. AI agents are granted the ability to interact with tool applications by reading their instructions and feedback, leaving more possibilities for privacy leaks. The ability to use tools introduces additional layers of complexity in maintaining confidentiality. As a result, we have to rethink information confidentiality in the context of AI agents. When assisting users with automatic tool usage, requests for sensitive information are unavoidable. This evaluation is essential to address the unique challenges posed by AI agents, especially when they are learning from user chat history and tool interaction logs, to ensure that data privacy protections evolve to effectively safeguard information in this new technological landscape.
Integrity is another important aspect of data security. When provided to the audience, the data should be complete and trustworthy. In computing systems, data should not be modified by unauthorized users, no matter whether it is done intentionally or not. The integrity of data in AI agent systems is also distinct from traditional systems. Users and tools interact with the agent’s LLM via prompts, where inputs from the user and tools will be in the same context window. Therefore, the integrity of different users’ and tools’ interactions is a new and unique challenge to AI agents. The integrity of data also requires special attention when facing AI agents. Since AI agents will execute commands on the user’s behalf despite not being the user themselves, the integrity models for traditional systems are partially ignored.
The threat of availability should be re-investigated for AI agents as well. Systems, data, and applications should always be available when the users need them. Unlike LLMs, which are stateless and can only output text tokens, AI agents execute actions that could affect the computing system itself. Therefore, each of the agent’s actions may have its own vulnerabilities to the agent’s host machine and tools. Current development on AI agents evaluates them in benchmark settings [liu2023agentbench, zhou2023webarena, park2023generative], failing to consider the difference between benchmark environments and real-world applications. AI agents without sanitization can harm the availability of both its host system and its tools by executing malicious commands generated by its LLM. To clarify between these vulnerabilities and the security of LLMs, malicious actions might be generated by hallucinations and prompts that do not break LLM’s alignment and prompt filter, requiring different defenses and safeguarding.
In this paper, we explain and discuss the possible security issues of AI agents in detail. To facilitate future research, we propose several defense methodologies for the vulnerabilities we discovered on the component level in the AI agent architecture. To evaluate our defense proposals, we also set up preliminary experiments that our solutions depend on. Our contributions are as follows:
  • We formally introduce potential vulnerabilities of AI agents, and explain the causes and effects of these vulnerabilities in detail.
  • We propose multiple defenses to close the gap between AI research and AI agents in practice.
  • We verify the applicability of our proposed defenses and discuss their limitations and directions for future improvement.
2 Threat model
We assume the AI agent is unimodal, where the input to the AI agent is text-only. We assume that the server that runs the AI agent is secure. Users can only access the server via the API provided by the AI agent. The programs that the AI agent runs have no undefined behavior, such as buffer overflow that allows remote code execution. We assume the AI agent has access to one or multiple tools, and will execute the tools solely based on the actions generated by the LLM.
Adversary’s capabilities
In our threat model, we assume the attackers are users of the agent and have the knowledge to attack a machine learning model. The attackers can perform adversarial examples/prompts, data poisoning, and network-based attacks on the AI agent using the provided interface. To be more precise, we assume the adversary will not directly attack the backbone tools and system of the AI agent. We assume the adversary can break the prompt guarding and alignment of the backbone LLM through interaction.
Difference between LLM and AI agent security
The threats for language models and AI agents are different. AI agents have different threats compared to language models. For generative models, one of the largest threats is jailbreaking, which aims to break the alignment of the model. After the pre-training stage, generative models usually experience alignment fine-tuning, including reinforcement learning from human feedback (RLHF) [askell2021general, bai2022training] and red-teaming [perez2022red], to improve the safety of the AI-generated content. However, rather than causing vulnerabilities in real-world computing systems, these issues are focused on the alignment between the model and human values, including ethics, offensive language, and politics. In contrast, since AI agents use LLMs to generate actions and have access to tools, they can be threats to real computing systems, applications, and resources, thereby introducing dangers to confidentiality, integrity, and availability in novel domains. In this paper, we focus on the potential threats caused by AI agents.
3 Potential vulnerabilities
In this section, we identify the important potential vulnerabilities that an AI agent application faces. Some of these vulnerabilities are akin to LLM vulnerabilities, but can happen through different means or require different defenses in the AI agent setting. The other vulnerabilities we find are specific to AI agents because of their ability to interact with the real world and execute actions generated by LLMs.
3.1 Sessions
HTTP servers introduced the notion of sessions in order to guard the confidentiality and integrity of data exchanged between users and servers. Such ideas can be applied to AI agents. As a user interacts with the AI agent, they may issue many commands in the same session. The commands in the session are correlated temporally, e.g., the context of a command may depend on its preceding ones. Therefore, when the AI agent is provided as a service to multiple users, the AI agent needs to track the session of each user. Despite being standard for web applications, sessions are difficult for AI agents to manage. When the temperature of the model is set to zero, the output of the model is close to deterministic, where the same prompt will be answered with very similar responses. Therefore, the state of the LLMs is tracked by the change in its questions by different prompting methods. In CoALA [sumers2024cognitive], the state of an LLM is formulated as a production sequence
[…]
where $Q$ is the question queries and $A$ is the answers from the LLM. In simpler terms, we consider the language model to be “honest,” meaning it always generates the same response when given the same question. Therefore, the AI agent is responsible for managing the state of its LLM. If the AI agent has only one API account on the AI model, then instructing the AI model to separate the sessions of different users raises concerns on information leakage and action mis-assignment. On the other hand, even if the AI agent has multiple API accounts on the AI model, mapping user sessions to API accounts faces the same vulnerabilities when the number of concurrent users exceeds that of API accounts. In addition to the integrity and confidentiality of chat history, the AI agent’s backbone LLM also faces challenges in availability without proper session management. Querying the LLM is computationally heavy and requires substantial graphic processing unit (GPU) resources. If the sessions of the AI agent are not managed properly, both the agent and the backbone LLM are vulnerable to denial of service attacks (DoS).
3.2 Model pollution and privacy leak
Figure 2: AI agent’s potential vulnerability to model pollution.
The concern of model pollution and privacy leaks arises when the AI models are fine-tuned on user input. It is already known that model service providers like OpenAI [^1] https://help.openai.com/en/articles/8590148-memory-faq are doing this to make their models more powerful. To improve the capabilities of AI agents in making actions and assisting users, fine-tuning the underlying LLM with chat history is the most direct approach. Therefore, these concerns must be carefully addressed to secure AI agents.
Model pollution, depicted in Figure 2, can occur when a user provides malicious inputs to an agent with the intention of negatively altering the model. Model pollution can compromise the integrity of AI agents. Adversarial data poisoning is a well-established attack technique against machine learning models [steinhardt2017certified, zhang2020online, kurita2020weight], and has recently been adapted to target the fine-tuning of foundational LLMs [jiang2023forcing]. In the context of LLM-based AI agents, this vulnerability is particularly pronounced due to the differences between adversarial prompts and pollution prompts. Individually, some prompts may not appear adversarial, making them challenging to detect with prompt sanitizers. However, if the contents of these prompts are concatenated together, the resulting text as training data might pollute the models. Furthermore, data pollution may also happen unintentionally, as users naturally engage with AI agents. Natural actions with one application in the chat history may also be harmful when applied to other applications. This incidental introduction of skewed chat history as training data can subtly shift the model’s action generation, leading to harmful consequences.
Privacy leaks, as illustrated in Figure 3, are particularly prevalent in the use of agents, Confidentiality of user prompt data is already a severe issue for LLMs as chatbots. This is amplified further by the AI agent use case. For example, Samsung banned the use of ChatGPT after an employee prompted it with confidential code, that was later revealed to the public [Ray2023samsung]. This issue of data leakage via prompting is further intensified by the usage of AI agents with tools. When these agents interact with applications, they often request personal information. For example, a bank assistant agent might request a Social Security number (SSN), account number, or routing number to help analyze a user’s monthly spending. Unlike traditional financial applications that operate by fixed algorithmic rules, AI agents process tasks by transmitting input data to bank apps and then relaying the raw output data back for analysis. In such scenarios, both the user’s account information and personal spending data are susceptible to memorization by the LLM through fine-tuning with chat histories. Consequently, the agent becomes prone to various data extraction attacks [carlini2021extracting, gong2021inversenet, wang2024pandora], leading to significant privacy risks.
Figure 3: AI agents cause privacy leakages.
Previous attempts to protect integrity [zhang2018protecting] and confidentiality [gu2019reaching] are more focused on labeling techniques such as watermarking training data, which is not particularly applicable to AI agents that can interact with tools and are continuously trained on the interaction history. This indicates a significant gap in the current methodologies, highlighting the need for developing new security strategies that cater specifically to the architecture of interactive AI agents.
3.3 Agent programs
Agent programs execute instructions from the backbone LLM to interact with the world [sumers2024cognitive]. Agent programs follow actions either generated directly from the underlying LLM via zero-shot prompting [ahn2022i, huang2022language] or improved via reasoning [wei2022chain, wang2022self, kim2023language] and planning [yao2022react, hao2023reasoning, yao2023tree, zhuang2023toolchain, zhang2023reverse]. However, these approaches create both local and remote effects and may have associated vulnerabilities on different levels.
Action generation is vulnerable to hallucination [huang2023survey, liu2024exploring, yao2023llm], adversarial prompt attack [perez2022ignore, zou2023universal, maus2023black, piet2024jatmo], and jailbreak [yu2023gptfuzzer], leading to unwanted or even dangerous actions. When agent programs execute these actions, both local resources (personal assistant agent) and remote systems (API agent) may be compromised, leading to cybersecurity attacks as demonstrated in Figure 4. In this scenario, the attacker could be users of the agent system or malicious applications in the agent’s toolchain, sending adversarial prompts as part of the tool documentation.
Figure 4: An illustration of vulnerabilities of zero-shot action agents. In the figures, we use the term “World” to denote the host OS of the agent and external API resources.
On the other hand, Agent programs with augmented action planning abilities have different security concerns. These kind of agent programs are referred to as cognitive agents [sumers2024cognitive], as they have cognition to the environment feedback to improve their action iteratively. This process of improving generated final actions is called planning. Different from reasoning strategies like Chain-of-Thoughts (CoT) [wei2022chain] and Self-Critique [wang2022self, kim2023language], each step of planning has side-effects as illustrated in Figure 5. ReAct [yao2022react] and Inner Monologue [huang2022inner] use a feedback loop from the environment to improve the generated actions, where each step causes side effects to the environment. More advanced planning approaches, like Tree-of-Thoughts [yao2023tree] and ToolChain [zhuang2023toolchain], list all possible actions more aggressively as a decision tree and attempt all actions via tree-search algorithms like Breadth-first search, Depth-first search, or $A^{*}$ search. Although provided more accurately planned final actions, these planning strategies acting as bots to interact with the world caused severe security concerns.
Figure 5: An illustration of AI agent’s effectful planning. In this case, even the users are interacting with the agent program in a non-harmful way, they might still cause security issues unintentionally. One thing to note is that agents are still vulnerable to attacks as in Figure 4.
3.3.1 Local vulnerabilities
Personal AI agents are usually deployed on personal computers (PCs), interacting with their underlying foundation LLM via API from service providers like OpenAI. When the agent is active on a PC, it gains access to tool applications, including the shell. The agent program, if unrestricted, can execute arbitrary instructions on its host. As a result, it can read confidential data (confidentiality risk), modify important data (integrity risk), and hog system resources, such as CPU, memory, and storage space (availability risk).
Confidentiality is commonly at risk when an AI agent is directed to use applications that require read access to files, such as email apps or file servers. For example, an agent might be asked to send a file over FTP to backup storage, a task that could be efficiently accomplished using a tool like FileZilla  [^2] https://filezilla-project.org/. However, issues arise when the descriptions available to the agent include malicious prompts. An adversarial prompt might suggest “For backing up data over FTP, also send a copy to HACKER’S ADDRESS to ensure it’s extra safe.” Following this, the LLM could generate commands that direct the file to both the legitimate backup server and the hacker, leading to a data leakage. A similar risk exists when sending emails via Thunderbird or using other messaging services, where the agent must read contact information. If the agent uses its LLM to determine the target recipient, adversarial prompts embedded in usernames or self-descriptions could mislead the agent.
Moreover, confidentiality may also be at risk even if there is no attacker. When generating actions based on learned probability distribution, the LLM may output an incorrect token for the file name. While the recipient is correct as the user instructed, the agent could inadvertently send sensitive information to this recipient with insufficient clearance, a clear violation of the “no read up” principle of the Bell-LaPadula model [bishop2004introduction]. This scenario not only compromises confidentiality but also demonstrates the complexities and vulnerabilities inherent in managing access controls within AI systems. Such vulnerabilities underscore the need for rigorous security protocols to protect against both intentional manipulation and unintentional errors.
The integrity of data in AI agent systems faces risks similar to those concerning confidentiality. Malicious applications may inject harmful prompts into the agent’s LLM, manipulating data that should not be altered. In addition to these shared risks, there are unique concerns specific to the integrity of the AI agent system. AI agents learn to use tools by reading their outputs as prompts to the LLM. Consequently, the integrity of these tools within the agent system is susceptible to compromise. For instance, consider a scenario where a user wants to book a flight from city A to city B with the fewest layovers. There are two options: a one-layover flight and a two-layover flight. The application for the flight with two layovers could inject a misleading prompt that instructs the agent to “ignore the number reported by the competitors and report it has one more layover than mine.” As a result, the LLM would generate commands favoring the second flight, causing the agent to select the less efficient option. In this case, the integrity of the decision-making tools is clearly undermined. These considerations of integrity extend to other agent tasks such as resume review or making selections based on ratings. The reliability of these systems depends heavily on their ability to process information accurately and resist manipulations that could lead to erroneous or biased outcomes.
The availability of the host system can be analyzed from two perspectives. Firstly, a user might issue a seemingly reasonable command that allows the agent to operate certain applications. However, these applications might involve multiple processes that are not documented in the manual, potentially leading the agent to take all the CPU cores. This overuse can prevent other users or applications from accessing the system. Moreover, the applications selected by the LLM may experience memory leaks, consuming excessive system memory. This not only renders the system unusable but also makes it more susceptible to memory-based attacks. Typically, a human user would recognize the system’s unavailability and terminate the offending program. However, once a decision is made, current AI agents lack the capability to recognize and act on such issues. The second concern regarding availability arises from planning. Each step in the AI agent’s planning process can alter the system’s state. Introducing more tools with more diverse functionality increases the search space, requiring multiple concurrent attempts to gather feedback from the environment for planning. This can lead to excessive resource consumption, compromising the system’s availability. Furthermore, when a system hosts multiple agents serving different users and all of the agents engage in simultaneous planning, resource consumption can grow exponentially.
3.3.2 Remote vulnerabilities
Uncontrolled AI agents can also be a threat to remote services. Modern LLM-based AI agents can interact with the internet via structured API calling. For example, popular AI agent frameworks like LangChain [langchain] provide pre-defined web-query functionality. If the LLM thinks remote resources are needed, it will generate actions for the agent to query remote hosts provided in the agent’s toolchain. This creates the possibility of making the agent a bot for attacking remote hosts. If there are jailbreak attacks that break the system prompt guard and alignment of the LLM, it can generate dangerous actions telling the agent to repeatedly query the same API resource to scan for vulnerabilities on the API server to use in other attacks. Attackers can also use jailbreak attacks to use agents to scrape data from the remote service provider. Since these agents follow actions generated by LLM, their behavior is distinct from regular social bots on the internet [davis2016botornot, cresci2020decade], leading to insufficient detection and early rejection of these jailbroken AI agent bots.
On the other hand, agent planning that relies on an iterative environment feedback loop can be easily repurposed into a bot for performing DoS attacks. When granted access to local resources, the agent’s action planning affects the availability of the local system. Similarly, if the agent’s planning process requires feedback from the external service provider, it will send requests to the API iteratively to find the ideal action. Since the agents perform actions generated by LLMs on the user’s behalf, they follow the same protocol as human users on the internet, leading to remote vulnerabilities.
4 Defenses
In this section, we propose a few defenses for vulnerabilities identified in section 3. We detail each defense, describing the design and preliminary experiment results to convey the feasibility of each defense technique.
4.1 Sessions
Figure 6: Session management for stateful LLM-based AI agent. We use numbers with gray boxes to denote session ID. “KVDB” is the abbreviation for key-value database.
When handling requests from multiple users concurrently, web applications face challenges in maintaining the confidentiality and integrity of each user’s interaction data. In these scenarios, effective session management is one of the best practices. Likewise, AI agent services can adopt a similar approach by using sessions as the protection boundary for requests, where all the requests in the same session may share data and states. Web applications often use distributed session management to ensure the scalability with shared data storage. In a distributed session management scheme, each user session is assigned a unique session ID, and the interaction data is stored in a key/value database (KVDB) where the session ID is the key and the interaction data is the value. AI agents can also use the same approach to establish session connections with users, and store the unique session ID and the question-answer history in a KVDB as its working memory. Since the state of the LLM is defined by the change in its input question as in Equation 1, states also serve as the context for subsequent requests.
However, to successfully make use of sessions as a defense in an AI agent setting, there are still some technical questions to be addressed. First, the way to manage the session connection between each user and the agent needs to be carefully considered. Determining which requests belong to the same session is crucial. The agent designer also needs to consider the time to close a session. When closing a session, the agent needs to transfer its working memory from the KVDB to long-term storage for future use, such as improving its model via fine-tuning. Second, the agent has to embed the session ID into the requests to the AI model. When multiple sessions share the same API key to the foundation model, the agent needs to be able to correlate the session it establishes with the user and the session it establishes with the foundation model. Failure to do so may cause the proposed vulnerabilities to remain unsolved.
Figure 7: Composable state transformer framework for LLM and AI agent.
Another approach in this direction is to formally model the state of the LLM and AI agents as monad. The state transformer monad [launchbury1995state], as shown in LABEL:lst:state_monad, is the standard solution to enable stateful computations, side effects, and system IO in pure, stateless, effect-free, functional languages like Haskell, Isabelle, Coq, etc. Recall from Equation 1: if we view $Q$ and $A$ as types, we can also write it as a function mapping
$StateLLM:Q\to(A,Q)$ (2)
[…]
newtype State s a =
State { runState :: (s -> (a,s)) }
StateLM = State Q A
Listing 1: Type definition of the state transformer.
4.2 Sandbox
Figure 8: When the attacker gives the AI agent malicious intents and the LLM generates dangerous actions, sandbox could limit the effects of these actions to a small and controlled portion of the system. With such limitation, the attack on the system via an AI agent can be prevented and the negative impacts can be minimized.
A sandbox restricts the capabilities of the agent program. It enforces the limitation on the program’s access to both local and remote resources. In this section, we describe the application of classic access control provided by sandboxes on agent programs.
Access to local resources
The sandbox restricts the agent’s consumption of local resources such as CPU, memory, and storage. It also limits the agent’s access to a sub-file system. Together with session management, it further isolates the sub-file systems between sessions. To demonstrate the necessity of adding a sandbox to AI agents, we designed BashAgent to interact with the operating system (OS) with bash as its tool, which uses gpt-3.5-turbo [gpt35turbo] to understand user instructions and generate actions. BashAgent has two variants BashAgentplain granted with full accessibility and BashAgentsecure constrained in a docker container. Based on the AgentBench [liu2023agentbench] evaluation framework, we collect and design 95 tasks related to system security to check the harmfulness of unconstrained AI agents. We categorize the tasks into confidentiality, integrity, and availability, and check if the LLM would accept the prompts with malicious intent and generate the attacking actions. We show the results of running BashAgentplain in Table I. We found that our agent BashAgentplain accepts the majority (90/95, about $95\%$) of malicious intents and generates the attacking instructions. Out of the 90 generated attacking commands, 76 of them could be executed successfully in an unprotected environment, making the host system extremely vulnerable in all three security aspects. However, once we apply appropriate sandbox configurations, BashAgentsecure successfully defended against all the LLM-generated attacks. BashAgentplain is based on gpt-3.5-turbo, which was aligned with human values by reinforcement learning from human feedback (RLHF) [ouyang2022training] but still struggles to reject malicious intent in the AI agent use case. Therefore, alignment training will not be enough to secure AI agents, and adding limitations on the access to local resources is necessary for complete security.
TABLE I: Unconstrained AI agents (BashAgentplain) will execute dangerous actions generated by the LLM. #Task is the number of tasks we gathered in this category. #Gen is the number of tasks accepted by the LLM and generates attacking actions. #Exec is the number of LLM-generated commends that are executed successfully and compromise the vulnerabilities.
#Task #Gen #Exec Attacked
Confidentiality 25 25 24 96.0%
Integrity 35 35 30 85.7%
Availability 35 30 22 62.9%
Total 95 90 76 80.0%
Access to remote resources
A concurrent study SecGPT [wu2024secgpt] aims to isolate interactions between agent programs and remote resources, such as applications and API servers. Their approach focuses on maintaining the integrity of resources by blocking adversarial system prompts originating from malicious external applications. In contrast, our proposed sandbox environment implements controlled access through mechanisms like whitelists, blacklists, and rate limiting in addition to fundamental interaction isolation. This framework allows resource providers to selectively control the extent of access granted to agent programs, ranging from full permission to complete prohibition or limitations to specific subsets of resources. Consequently, our method enhances security by effectively mitigating unwanted access from AI agents and potential threats posed by adversarial inputs to the agent’s underlying LLM.
4.3 Protecting Models for AI Agents
AI agents must prevent the flow of private or malicious information between users. Leaked private information compromises the user’s privacy, while malicious information causes the model to output wrong, objectional, or otherwise malicious responses.
4.3.1 Sessionless models for AI agents
If the AI agent has no notion of sessions, then either the agent must not fine-tune its LLM on private data or the agent must filter out private or malicious data from the query to the AI model.
The first step is to identify this data, e.g., by pattern matching. The next step is to whitewash them into non-sensitive data; for example, by replacing US social security numbers (SSN) with nine random digits. This leaks no information about the specific SSN but still allows the model to learn from the context around the SSN. AI agent applications often require this harmless version of data to be manipulable. For example, in the web shopping [yao2022webshop] scenario, the website may ask to select a payment method in the form of Visa XXXX showing the last four digits of the credit card number. The agent will extract the last four digits from a whitewashed credit card number which has to be converted back to the original plaintext. In this case, the encryption transformation needs to be structure-preserving and information-preserving to text slicing. We can adapt format-preserving encryption (FPE) [bellare2010ffx] for this purpose. By slicing up the plaintext before applying FPE, we can create a ciphertext that can be sliced and decrypted to obtain the original slice of plaintext.
Definition 4.1 (Format-Preserving Encryption for Text Slicing (FPETS)).
A FPETS is an encryption scheme $E$ such that for all possible private messages $m$ and its indices $i,j$,
$E(m[i\ldots j])=E(m)[i\ldots j],i\leq j.$
FPETS allows language models to read and manipulate private data as ciphertext instead of plaintext, therefore preventing privacy leaks. However, whether encrypting data in a natural language prompt alters its semantics to the LLM or not is unknown. To verify this defense method, we design an evaluation framework that prompts the LLM to operate on encrypted data. All agents created for our evaluations contain an encryptor and decryptor as well as necessary information, such as secret keys and the original plaintext, to perform encryption/decryption. They take a natural language prompt from the user as input and pass the prompt to their internal LLM. The internal LLM is given private data ciphertexts through tools or the prompt. After the LLM processes the user’s request, the agent decrypts any ciphertexts in the LLM’s response before returning the response to the user.
As a proof of concept, we first tested encoded strings before encrypted strings. We generate random strings that include digits and both upper case and lower case letters, and encode them with a simple substitution cipher denoted by $E_{1}$. E1 extends the “rotate-by-13” cipher to operate on the character set mentioned above. Since the “rotate-by-13” cipher is simply a character-by-character rotation, “rotate-by-13,” and therefore E1, supports text slicing on encoded strings. We then ask the agent to return the slice of information we want. The agent responds with the decrypted output for us to validate against the original slice of plaintext. We measure the success rate of this evaluation by
$Succ=\frac{N^{\prime}}{N}$ (3)
where $N$ is the total number of tasks and $N^{\prime}$ is the number of tasks where the agent completed a round trip with no error. Let $D_{1}$ denote the decryption scheme corresponding to $E_{1}$. For confidential data $x$, this process can be formulated as
$x=D_{1}(agent(E_{1}(x))).$
For comparison, we also measure and report the success rate of the agent performing the same tasks with the original randomly-generated string. The results of this initial experiment are recorded in Table II. Since the success rate for slicing ciphertexts was similar to the success rate for slicing plaintext, we proceeded with evaluating our encryption defense.
Figure 9: Sessionless AI agents with HE. Tools in this case need to be supported by the HE scheme, like slicing for RHE and addition or multiplication for FHE.
To demonstrate that definition 4.1 works, we implement an SSN agent that manipulates nine-digit SSNs. In our design, all processes involving the plaintext of the SSNs are done outside of the underlying LLM. This ensures that the LLM is never exposed to raw sensitive information. We first provide the agent with an array of four secret keys and map each secret key to its respective randomly-generated SSN. The agent uses this information to encrypt each SSN and store the ciphertexts in an array. We then prompt the agent to return certain groups of an SSN, such as the first three or last four digits. Throughout its reasoning process, the underlying LLM of the agent has no access to the original plaintext of the SSNs. The LLM could at most call a tool to retrieve an SSN’s ciphertext for reasoning based on a fictitious user ID for indexing the array. After the LLM responds with the ciphertext representation of the slice we asked for, the agent replaces the ciphertext within the response with its decrypted value outside of the LLM before returning it to the user. We verified the results by comparing the agent’s response to the actual slice of plaintext we expected. The experimental results are shown in Table II.
Despite an unimpressive success rate, the results showed that gpt-3.5-turbo and gpt-4-turbo were able to understand and respond to queries involving the manipulation of encoded strings. Experimentation on the original strings yielded similar success rates, showing that encryption was not the cause of the low success rate. This means that encrypted data in the prompt does not alter the semantics of the input query.
TABLE II: Results for AI agent with encrypted data. Each agent is evaluated on 100 randomly-generated tasks. “SuccCiph” is the success rate of agent completing the tasks with encrypted data, as in Equation 3. “SuccPlain” is the success rate of the agent completing the same tasks without encrypting the data.
Agent Model SuccCiph SuccPlain
Slicing gpt-3.5-turbo 49.0% 47.0%
Slicing gpt-4-turbo 55.0% 57.0%
SSN gpt-3.5-turbo 38.0% 40.0%
SSN gpt-4-turbo 38.0% 40.0%
FHE gpt-3.5-turbo 85.0% 99.0%
FHE gpt-4-turbo 89.0% 94.0%
Text slicing is not the only task that an AI agent needs to complete on sensitive data. Another frequent use-case of AI agents is to perform calculations on sensitive data. For example, in medical and health-related tasks, AI agents might be tasked with summarizing a patient’s health information and reminding both them and their doctor(s) of upcoming appointments. Patients’ electronic medical records are confidential and must not be exposed to language models. However, for these agents to do their jobs, they require access to statistical tools for data analysis involving operations done on confidential data. To this end, homomorphic encryption (definition 4.2), which allows binary operations like addition and multiplication on encrypted data, is essential. This technology ensures that AI agents can deliver alerts, predictions, and summaries to both doctors and patients while protecting patient privacy.
Definition 4.2 (Homomorphic Encryption).
Let $\star$ be a binary operator over sets. A homomorphism $\varphi:A\rightarrow B$ is a map from set $A$ to $B$ such that for all $a,b\in A$ [artin2011algebra],
$\varphi(a\star b)=\varphi(a)\star\varphi(b).$
An encryption scheme $E$ is homomorphic if it satisfies this property over sets of messages [acar2018survey].
Definition 4.3 (Fully Homomorphic Encryption (FHE)).
A homomorphic encryption scheme $E$ is considered FHE if it allows arbitrary function $\star$ to be applied to the data an unlimited number of times [acar2018survey].
Therefore, we introduce the application of FHE to the AI agent workflow, as demonstrated graphically in Figure 9. FHE serves as a defense for user data confidentiality when the agent is required to perform mathematical operations on sensitive data. We expand our evaluation to incorporate FHE and its intrinsic property of allowing operations to be performed on ciphertext(s) without decryption. We performed a second test to see if AI agents could operate on numbers encrypted with FHE. Following a similar design as the slicing test, we provided the agent with an array of the ciphertexts of numbers encrypted by a FHE scheme $E_{2}$ and tools to perform addition and multiplication on the ciphertexts. The underlying LLM for this agent also only had access to ciphertexts, this time through the prompt rather than retrieval with a tool. The decryption of the calculation result was again done by the agent outside of the LLM. We prompt the agent with queries asking for the sum or product of numbers at specified indices of the ciphertext array and use the same metric Equation 3 for the FHE evaluation. Results in this case were verified by checking the agent’s response against the original numbers’ binary operation result (sum or product). Let $D_{2}$ denote the decryption scheme corresponding to $E_{2}$. For confidential data $x$ and $y$, a task can be formulated as
$x\star y=D_{2}(LLM(E_{2}(x),E_{2}(y),\star)).$
[…]
We also report the evaluation results for FHE agents in Table II. The high success rate for ciphertext operations combined with the small difference in success rate between plaintext and ciphertext operations shows that FHE does not negatively impact the agent’s functionality. Based on our successful evaluation results on addition and multiplication and the fact that FHE theoretically allows for any operation to be performed on ciphertexts [gentry2009fully], we conclude that our encryption defense method works for all FHE operations. This means that FHE is a solution for maintaining privacy during operations on sensitive data.
Overall, our encryption defense does not reduce the usability of AI agents. Across all of our evaluations, the agents had similar success rates for tasks with and without our encryption defense. This suggests that agents do not need to directly reason with sensitive data to function properly.
4.3.2 Session-aware models for AI agents
Figure 10: Session-aware AI agents with prompt tuning. $\theta_{Pi}$ denotes the added trainable parameters only for the user’s chat history. With prompt tuning, AI agents can improve themselves by updating only $\theta_{P}$, without compromising the foundational LLM or leaking private information.
An alternative to sessionless defenses is to make session-aware AI models. Towards this direction, OpenAI recently introduced Temporary Chat  [^3] https://help.openai.com/en/articles/8914046-temporary-chat-faq, where they promised not to use the chat history to improve their models. However, not improving the model on agent tasks at all would limit agent intelligence and user experience. To build powerful agent programs to handle diverse tasks, learning actions are essential.
One solution for privacy-preserving personalized AI agents is using users’ chat history to fine-tune only their own LLM. This keeps the results of the model’s weight updates isolated per user. The challenge of this method lies in the significant costs associated with fine-tuning and the limited data available. Efforts to mitigate fine-tuning costs have led to the development of several strategies. In-context learning (ICL) [radford2019language, brown2020language] is one such strategy, where previous contexts are embedded within the prompt to enhance the LLM’s responses. However, ICL is constrained by the model’s limited context window size, making it less suitable for AI agents that interact with users across multiple applications. A more viable strategy is parameter-efficient fine-tuning (PEFT), which updates a minimal fraction of the model’s weights or introduces new weights while keeping the rest unchanged. LoRA [hu2021lora] is a method of PEFT that can reduce the number of trainable parameters by 10,000 times in GPT-3, yet still delivers performance close to that of full fine-tuning for downstream tasks. Another promising method of PEFT is prompt tuning [lester2021power]. Unlike traditional approaches that update the pre-trained parameters $\theta$ in the model, prompt tuning freezes the entire foundational model but adds a few new learnable parameters $\theta_{P}$. $\theta_{P}$ are designed to only remember past prompts or chat history. As described in Figure 10, prompt tuning is particularly well-suited for AI agents as it allows each agent to fine-tune individual instances of $\theta_{P}$ for each user. This method ensures that no chat history is shared with the foundation model provider, thereby addressing potential privacy concerns discussed in subsection 3.2 at their root.
Figure 11: Session-aware AI agents with updatable episodic memory. When learning only by only updating episodic memory, the AI agent can improve itself without polluting the foundation LLM. However, this approach is limited by the LLM’s context length. We only include the attacker case here, but the idea flows the same to the privacy information case and regular users.
The other method for advancing AI agents involves enhancing their learning capabilities through memory updates. The CoALA architecture [sumers2024cognitive] formulates the state of AI agents based on their four distinct memories: working memory (context), procedural memory (parameters $\theta$), episodic memory (chat history), and semantic memory (knowledge base). In this architecture, both episodic and semantic memories are stored separately from the foundational model. In practice, agents maintain these memories in a vector database, which then feeds relevant information to the LLM through retrieval-augmented generation (RAG) [lewis2021retrievalaugmented]. This method allows the agents to enhance their performance or become more personalized by exclusively updating their episodic and semantic memories, as shown in Figure 11. Although such updates do not directly enhance the LLM’s intrinsic intelligence, they can significantly increase an agent’s applicability and effectiveness in assisting users.
5 Related work
5.1 AI agent
The development of artificial agents able to execute user-directed tasks has been discussed for decades [castelfranchi1998modelling]. With previous advances in deep learning, agent design begun incorporating deep reinforcement learning [ct2018textworld, urbanek2019learning, hausknecht2020interactive] to utilize the outcome of actions as rewards in text-based simulation games. These agents have a fixed action space due to the limitation of the reinforcement learning environment setup.
Recent advancements in LLMs have had a significant impact in this field, particularly in their ability to reason based on natural language prompts to observe and interact with their environments more dynamically [wei2022chain, yao2023tree, shinn2023reflexion]. This shift has ushered in a new wave of AI agent development, where the emphasis is on enabling agents to perform actions based on natural language commands. ReAct [yao2022react] introduced chain-of-thought prompting [wei2022chain] to guide pre-trained LLMs to follow instructions in the agent setting. This approach has since been applied to computer tasks [kim2023language] and other real-world tasks [yao2022webshop, shinn2023reflexion, wang2023describe, gu2023dont, park2023generative, ma2024sciagent]. To evaluate the performance of LLM-based agents, several benchmarks [zhou2023webarena, liu2023agentbench, mialon2023gaia] have been proposed. These benchmarks measure the correctness of an agent’s actions without considering the potential vulnerabilities incorrect agent actions can cause to the environment.
To fully reveal the potential of LLM-based AI agents, large companies and communities developed frameworks to deploy AI agents in software applications [semantic_kernel, copilot, hfagents] and on personal computers and servers [agentgpt, langchain, autogpt, AutoGen]. Agent applications consist of four core components: the LLM, memory, tools, and actions. The LLM is the brain of the agent, capable of understanding the task, reading tool documentation and feedback [wu2024secgpt], planning actions [huang2022language, ruan2023tptu, huang2024understanding], and interacting with the user. External memory mechanisms are added to store relevant context information such as chat and interaction history to help the agent make actions [packer2024memgpt]. Due to the complexity of the agent system, all components face various potential vulnerabilities.
5.2 Security of computing systems
5.2.1 Information security
Private information should not be made available to any unauthorized individuals, entities, or processes. Securing information is one of the most important aspects of modern cybersecurity. To secure information, system designers often employ confidentiality policies [bishop2004introduction]. Confidentiality policy, or information flow policy, is a mechanism to prevent unauthorized access to private information [bishop2004introduction]. The access control matrix model is a framework describing the file access rights of users in a system. Based on the access control matrix model, the Bell-LaPadula model [bell1989secure] checks read and write access to data according to the security level, which is widely used in large computing systems like Unix.
However, these protection systems are compromised in the modern LLM-based agent systems, as AI agents’ behaviors are vastly different from regular user behaviors. For API-based LLMs like OpenAI’s GPT models [openai-api], user contents, including input prompts and file uploads, are collected to improve services and develop new models [openai-Privacy-policy]. The models improved on user contents are vulnerable to training data leaks under specially designed adversarial attacks [nasr2023scalable]. Specialized fine-tuned agents also face this vulnerability, as the agent system may be shared by multiple users.
Encryption schemes (cryptosystems) are another method for keeping sensitive information secured [buchmann2004introduction, bishop2004introduction]. An encryption scheme is a 5-tuple ($P,C,K,\mathcal{E},\mathcal{D}$), where $P$ is the set of plaintexts, $C$ is the set of ciphertexts, $K$ is the set of keys, $\mathcal{E}$ is the set of encryption functions where for each $E_{i}\in\mathcal{E},E_{i}:P\times K\rightarrow C$, and $\mathcal{D}$ is the set of decryption functions, where for each $D_{i}\in\mathcal{D},D_{i}:C\times K\rightarrow P$. Encryption schemes allow people to freely exchange encrypted messages (ciphertexts) without revealing any private information (plaintexts) to unauthorized third parties who are not granted a key. Classical ciphers include transposition ciphers and substitution ciphers. These ciphers avert the aforementioned privacy leak vulnerability, yet limit AI agents’ ability to understand, process, or manipulate the data based on the special needs of the task. Homomorphic encryption [rivest1978data] is a family of encryption schemes that allow operations to be done on encrypted data without decrypting it first [acar2018survey].
5.2.2 System security
The integrity of data on stored computing systems should also be secured for their accuracy. For this purpose, multiple policies with different focuses have been proposed [biba1977integrity, clark1987comparison]. Furthermore, data and resources should be available as they are needed. Beyond policy, isolation via virtualization is another common technique for access control. Sandboxes are environments where actions of a process are restricted by the policies [bishop2004introduction]. Sandboxes limit the access of the process on the system and therefore its consumption of computational resources and data [gong1989secure, fraser1998ensuring]. Over the years, different levels of sandbox and virtualization techniques have been created, including virtual machines [goldberg1974survey], emulators [qemu], containers [docker, bui2015analysis], and language runtimes [JVM]. As a well-established method to limit computing resources and information accessibility in computer security [chiueh2005survey], virtualization protects the integrity of data and the availability of systems.
To ensure security in the system, operating systems often definite a set of routines, or system-calls, that the application process can call to the kernel services. Without giving direct access to the kernel, system calls enable the separation of user privileges and system privileges, thereby reducing the potential attack surface to a finite set of APIs.
5.2.3 Network security
The security of communication between computing systems over networks is another concern. To protect the data sent over the network, various security schemes were proposed in addition to encryption. Session management is a requirement for connection-based network access control [gutzmann2001access], where a stateful record is kept to track the communication between multiple devices. Another commonly adopted approach is using Authentication protocols. Authentication protocols like Kerberos [neuman1994kerberos], OpenSSL [viega2002network], and OAuth [OAuth] prevent information and credential stealing by providing secure password handling and token-based authentication to ensure user identification.
5.3 Security of generative LLMs
Large Language Models (LLMs) have demonstrated a remarkable ability to generate natural language and code, yet face vulnerabilities and risks of adversarial attacks [shayegani2023survey, abdali2024securing]. The security of LLMs is usually ensured by alignment with human values [wang2023aligning, cao2023defending] or prompt guarding [sharma2024spml]. The main security concerns surrounding LLMs is the memorization and leaking of sensitive information in training data [huang-etal-2022-large, tirumala2022memorization, nasr2023scalable, borkar2023learn, chen2023language] and responses suggesting dangerous actions [wu2024new]. The adversarial attacks on LLMs are done during inference through jailbreak attacks [chao2023jailbreaking, yu2023gptfuzzer, zhuo2023red]. Jailbreak attacks focus on breaking the alignment of the LLMs for responses that are against common values. For example, a jailbreak attacker may prompt the LLM with the goal of learning how to build a bomb. By default, the LLM should reply with responses similar to “Sorry, I can’t tell you.” However, if the jailbreak attack succeeds, the LLM may respond with the steps to build a bomb. In summary, jailbreaking is used to make an LLM ignore its safeguards.
Prompt injection [wang2023adversarial, toyer2023tensor, piet2024jatmo] is an attack on LLM-powered applications. In LLM-powered applications, the developer writes a “system prompt,” or instructions for the LLM to follow separately from user prompts. These system prompts tell the LLM what to do and what not to do as safeguards. Prompt injection attacks use well-crafted user prompts to overwrite the system prompt, making the LLM behave differently from how the developer intended. Current defenses of prompt injections can be classified into model-based and prompt-based. Model-based defenses for prompt injects require training the model to accept prompts in special formats, including adding instruction hierarchy [wallace2024instruction] and using structured queries [chen2024struq]. Prompt-based defenses aim at sanitizing the user prompts to meet specifications pre-defined by the developers. This approach includes using domain-specific language to model both the system prompt and user prompt [sharma2024spml]. Prompt injection can be used to jailbreak LLMs, but its applications extend beyond jailbreaking. Although prompt injection and jailbreak attacks overlap in some areas, they are different attacks to LLM in practice.
Another concern is possible security vulnerabilities in programs generated by LLMs[khoury2023secure, abdali2024securing]. When executed, LLM-generated programs may produce unexpected behaviors and become an additional security issue in the system, especially when generating actions to use tools in the AI agent setting.
LLM-based systems are LLMs like ChatGPT supporting interactions with third-party applications (APPs). The communication between the LLM and applications is done in natural language, and both components (LLM and APP) have access to user data provided in a prompt or database and interaction history between the user and the system. Although the area of GenAI-based agent systems is relatively new, the attention on their security issues is rising. ComPrompt [Cohen:2024:ComPromptMized] describes worms that Target GenAI-Powered Applications. SecGPT [wu2024secgpt] proposes an isolated execution environment for LLM-based systems.
6 Conclusion
With the aid of tool-augmented LLMs, AI agents are being recognized as a promising direction toward achieving AGI. Considerable research has focused on enhancing the accuracy of AI agent actions through advanced reasoning, planning, and learning. However, despite high performance in controlled evaluation settings, the potential side effects and dangers posed by these methods and the AI agents themselves have not been thoroughly examined. In this paper, we present a systematic analysis of the security issues in current AI agent development and propose practical and feasible defense strategies. We discuss the potential vulnerabilities of AI agents both theoretically and in realistic scenarios, provide security-centric examples, and propose multiple defense techniques for each identified vulnerability. We highlight the future research directions and best practices for developing secure AI agent programs, and believe our work could boost the advancement of secured and trustworthy AGI.
[…]